Adv. Math. Econ. 14, 161-176 (2010)

Good locally maximal programs for the Robinson-Solow-Srinivasan (RSS) model

Alexander J. Zaslavski

Department of Mathematics, The Technion-Israel Institute of Technology, 32000 Haifa, Israel (ajzasl tx.technion.ac.il)

Received: July 30, 2008 Revised: May 29, 2009

JEL classification: D90, C61

Mathematical Subject Classification (2000): 49J99, 54E52

Abstract. In this paper we obtain results on the good choice of techniques in the long-run in the model proposed by Robinson, Solow and Srinivasan. We study this model with a nonconcave utility function which represents the preferences of the planner and establish some properties of locally maximal programs. We obtain a useful estimation for consumption over good locally maximal programs and show that a limit of good locally optimal programs is also a good locally maximal program.

Key words: Good program, locally maximal program, RSS model